### v1

# An agent-based model of behaviour change calibrated to reversal learning data

Abstract
Behaviour change lies at the heart of many observable collective phenomena such as the transmission and control of infectious diseases, adoption of public health policies, and migration of animals to new habitats. Representing the process of individual behaviour change in computer simulations of these phenomena remains an open challenge. Often, computational models use phenomenological implementations with limited support from behavioural data. Without a strong connection to observable quantities, such models have limited utility for simulating observed and counterfactual scenarios of emergent phenomena because they cannot be validated or calibrated. Here, we present a simple stochastic individual-based model of reversal learning that captures fundamental properties of individual behaviour change, namely, the capacity to learn based on accumulated reward signals, and the transient persistence of learned behaviour after rewards are removed or altered. The model has only two parameters, and we use approximate Bayesian computation to demonstrate that they are fully identifiable from empirical reversal learning time series data. Finally, we demonstrate how the model can be extended to account for the increased complexity of behavioural dynamics over longer time scales involving fluctuating stimuli. This work is a step towards the development and evaluation of fully identifiable individual-level behaviour change models that can function as validated submodels for complex simulations of collective behaviour change.
individual-based model, agent-based model, behavioural change, reversal learning, calibration, approximate Bayesian computation
I Introduction
Behavioural changes are fundamental to many real-world phenomena we seek to understand or predict through computer simulations. Such computational models describe how the actions of individuals in a system change in response to events caused by internal or external factors. Variation in behaviour over time affects the dynamics of the modelled real-world systems and the phenomena that emerge from them, from how infectious diseases spread in a population to how animals adapt to new habitats Verelst2016 ; Weston2018 ; Bauer2013 ; Leimar2024 . However, in many computer simulations, these individual-level submodels are based on strong assumptions about the underlying factors that influence behaviour and the conditions that drive behaviour change, with limited empirical support. Hence, such models cannot be validated or calibrated based on behavioural data, raising concerns about the verifiability of their results Verelst2016 ; Weston2018 . There remains a need for explicit computational representations of behaviour change that are supported by quantitative empirical evidence while maintaining the mechanistic adequacy to simulate observed and counterfactual scenarios. To help address this need, we present a simple (two-parameter) stochastic model of individual behaviour change and demonstrate that its parameters are identifiable from empirical time series data from reversal learning experiments.
To introduce our model, we begin with the general premise that individual decisions about behaviour map available information (stimuli) to a set of actions (responses). As a motivating example, consider endogenous protective behaviours during an infectious disease outbreak: The choice of some individuals to adopt social distancing practices depends on external information from their environment which could be globally available from (e.g.) government and media reports of disease statistics Collinson2015 ; Xiao2015 . Additional local information may also be available from communication and observations involving neighbours, friends, and family Andrews2015 . Such external information can support or contradict information internal to individuals, such as their perceived risks and benefits. How individuals then incorporate this information into their decision-making process may vary from person to person. Such variation often correlates with observable factors such as age, occupation, or socioeconomic status, which may reflect different capabilities and considerations that individuals have when making decisions Gozzi2021 ; Gauvin2021 ; Zachreson2021 . Thus, when modelling behaviour in computer simulations, design choices are often made about the types of information individuals can use to make decisions and the mechanisms by which they translate that information into actions (see, e.g., Durham2012 ).
In behaviour change models, environmental stimulus could be the number of infected people in an infectious disease model Poletti2011 ; Durham2012 ; Bayham2015 , the physical features of prey in an animal foraging model Dyer2014 ; Leimar2024 , or a reward indicating the desirability of a chosen action from the perspective of a global observer with complete knowledge Xia2021 ; Le2023 . In simulations, there are two general approaches to modelling how individuals act on this information: rule-based or learning-based. Rule-based approaches condition an individual’s behaviour on external information based on predefined threshold criteria or logical relationships Durham2012 ; Bayham2015 . In contrast, learning-based approaches assume that individuals are rational and behave optimally by choosing actions that maximize an objective function. The objective functions can be elaborate and may incorporate computing the costs and benefits of multiple strategies or examining past experiences Poletti2011 ; Dyer2014 .
All studies in which such models are developed must make assumptions on the factors that drive behaviour change, but many do so without being calibrated or validated from empirical data Verelst2016 ; Weston2018 . In those studies that use behavioural data, the information is often in the form of surveys which are costly to conduct and typically quantify behavioural intentions rather than actual behaviours. Additionally, most previous modelling studies do not draw from the rich body of literature that examines and describes the psychology behind behaviour change which has the potential to provide a stronger theoretical foundation for computational models Weston2020 . By designing a computational representation of behaviour that is informed by both experimental data and behavioural change theories, we can improve the validity, reliability, and practicality of computer simulations.
Figure 1: General setup for a reversal learning experiment. An individual selects between two options: A and B. Initially, one option is associated with a positive stimulus and the other option with no stimulus. After the reversal point, this reward association switches. The experiment tests the individual’s ability to adapt its behaviour to the reversal of the reward signal.
To develop our individual-based model of behaviour change, we aim to simulate the reversal learning experiment that is often conducted in neuroscience, psychology, and ecology to study how individuals adapt their behaviour to changes in their environment. An overview of the reversal learning experiment setup is shown in Figure 1. In a typical reversal learning experiment, an individual is tasked to choose between two options: a rewarded and a non-rewarded option. The experiment is repeated in a series of trials and examines whether the individual can learn which of the two options is rewarded. After some time, the reward association is reversed. The individual’s response to the change in the stimulus is then recorded over a second set of repeated trials over which the new reward association may be learned Rayburn-reeves2011 ; Dyer2014 ; Hassett2017 ; Xia2021 ; Le2023 ; Leimar2024 .
By simulating reversal learning, we developed an agent-based model of behaviour change with parameters that are fully identifiable from empirical data. We show that this model satisfies three characteristics of a good behavioural change model: (1) mechanistic adequacy, (2) tangibility, and (3) tractability. By mechanistic adequacy, we refer to the degree to which the model’s underlying processes are relatable to the features of the behaviour change process (in this case those characterising reversal learning). Tangibility means the model is tangible: its parameters have intuitive meaning and it produces output that can be directly compared to observable quantities or patterns. Tractability means the model is tractable: its parameters are identifiable based on comparison with experiments and it can be calibrated efficiently in practice with sufficiently low computational cost.
Below, we provide a detailed description of our agent-based model and reversal learning simulations (Methods, Section II). We then present the calibration and validation results (Results, Section III), followed by a discussion of the implications of our study (Discussion, Section IV). We conclude by giving a summary of our findings (Conclusion, Section V).
II Methods
In this section, we discuss how our agent-based model represents reversal learning behaviour. We then describe the experimental data we used to understand and evaluate this behaviour and the experiments that generated the data. Finally, we describe how calibration was performed.
II.1 Model description
We developed a simple agent-based model of behaviour change where an agent’s behaviour mimics that which is observed in reversal learning experiments. An overview of our agent-based model is shown in Figure 2. Abstractly, agents represent any individual capable of making decisions that could produce reward signals. In the model, the agent decides on an action $a$ from a given set of possible actions at every iteration of the simulation. We designed the agent to be capable of sensing its environment and remembering its past experiences which enable it to acquire and learn information for deciding which action to choose, resembling individuals in the real world. The information from the environment comes in the form of a stimulus $s$ that encapsulates the impact of the agent’s action on the environment in terms of an associated reward signal. The information internal to the agent is encoded as a memory value $m$ that quantifies the desirability of an action based on the agent’s past experiences. Given this information, the agent decides whether to stay with or change its current action based on the probability returned by its logistic decision function. The agent’s behavioural tendencies are characterized by the timescale of its memory $\tau$, and the logistic decision function’s steepness $k$ which specifies the agent’s certainty about the value of the information it has accumulated. In every simulation, the environment periodically switches the stimulus associated with the agent’s actions and observes the agent’s probability of choosing the rewarded or non-rewarded actions over time. Below, we provide a full model description following the ODD (Overview, Design concepts, Details) protocol for describing individual- and agent-based models Grimm2006 , as updated by Grimm2020 .
Figure 2: Schematic of the decision-making process in the agent-based model. (a) The agent chooses an action $a_{t}$ at time step $t$ and receives a corresponding stimulus $s_{t}$ from the environment. (b) In choosing the next action $a_{t+1}$, the agent considers the utility $u_{t}$ of its current action $a_{t}$. This utility is the sum of the instantaneous stimulus $s_{t}$ from the environment and the agent’s internal memory $m_{t}$ of past actions and stimuli with a timescale determined by the memory decay $\tau$. The agent decides the next action $a_{t+1}$ using a logistic function with a steepness $k$, which indicates the agent’s certainty on the informational value of any given utility.
II.1.1 Purpose and patterns
The model’s specific purpose is to illustrate a simple computational representation of a single agent’s behaviour that can be calibrated and validated from empirical data. Having such a representation of behaviour that is mechanistically adequate while being empirically identifiable will enable the incorporation of validated behavioural submodels into simulations of multi-agent systems. The patterns used as the criterion for evaluating the model’s suitability for its purpose are the time series of an agent’s decisions (for an individual) and the time series of the success rate (for an ensemble).
II.1.2 Entities, state variables, and scales
There are two entities in the model. The first entity is the environment which facilitates the reversal learning experiment. The state variables of the environment are the list of possible actions that the agent can choose from, the stimulus $s$ that is given to the agent for each action, the number of trial periods $n$, and the number of trials $r$ in each trial period. All the environment’s state variables can be specified from the setup of empirical experiments and are static variables except for the stimulus $s$ associated with each action which changes per trial period.
The second entity in the model is the agent whose behaviour is observed during the experiment. The state variables of the agent are its chosen action $a$ at every time step, the corresponding stimulus $s$ that it receives from the environment, a memory value $m$ which represents the agent’s preference for the chosen action $a$ based on its memory of past experiences (with timescale determined by the memory decay $\tau$), and a logistic function with steepness $k$ which represents the agent’s decision function. The action $a$, stimulus $s$, and memory $m$ are dynamic variables that change over time, while the memory decay $\tau$ and steepness $k$ are static parameters that characterize an agent’s behavioural tendencies.
Space is not represented in the model. Time is represented in discrete increments with an abstract unit of 1 trial and an extent equal to the number of trials $r$ multiplied by the number of trial periods $n$ in a given reversal learning experiment. This representation of time follows the setup used in the empirical experiments to which we refer for calibration and validation Buechel2018 ; Boussard2020 .
II.1.3 Process overview and scheduling
At every time step $t$, the model runs a simple 2-step process:
  1. 1.
    The agent chooses an action $a_{t}$ by executing the Choose_action submodel.
  2. 2.
    The environment gives the agent a stimulus $s_{t}$ associated with the agent’s chosen action $a_{t}$. The stimulus $s_{t}$ associated with each action changes per trial period.
II.1.4 Design concepts
The design concepts implemented in the agent-based model are basic principles, emergence, adaptation, learning, sensing, stochasticity, and observation. The agent-based model includes no objectives, prediction, and collectives.
Basic principles. The model is designed to demonstrate how behaviour could be represented in computer simulations such that it can be calibrated and validated from observable behavioural data. Moreover, the components we included in our computational model are informed broadly by the COM-B model of behaviour change which is a widely used theoretical framework on what drives behavioural changes (see Discussion for more information).
Emergence. The main result from the model which is the observed success rate over time emerges from the stimuli that the agent receives from the environment as well as its memory of past actions and stimuli. This probability of choosing an action is also partly imposed by the agent’s parameters for the memory decay $\tau$ and the steepness $k$ of the decision function which specifies the behavioural tendencies of the agent.
Adaptation. The agent has one adaptive behavior which is to choose an action at every time step. The agent’s decision of which action to choose is a probabilistic process where the probabilities are determined by a logistic function that takes as input the sum of the instantaneous stimulus that the agent receives from the environment and the agent’s internal memory of past experiences.
Learning. The adaptive behaviour of the agent incorporates learning through the use of an internal memory that stores the agent’s past actions and their associated stimuli. This memory enables the agent to learn the association between actions and stimuli and is developed as the agent gets more experience. The influence of past actions and stimuli on the agent’s decision at every time step decays exponentially over time. This learning process is detailed in the Get_memory submodel (described in Section 2).
Sensing. The agent is assumed to sense exactly the instantaneous stimulus associated with its action (it knows whether or not it made the rewarded decision). This sensing mechanism provides the agent with immediate feedback on the desirability of its most recent action.
Stochasticity. The action that the agent chooses at each time step is designed as a stochastic process to represent non-deterministic behaviour.
Observation. The fraction of agents choosing the rewarded action over time is observed from the relative frequency with which the agent chooses the action across an ensemble of simulation runs.
II.1.5 Initialization
At initialization, the list of possible actions that the agent can choose from and their associated stimulus are specified. Here, the agent can choose between options A and B. One of these options is initially assigned with a positive stimulus of +1+1, while the other has a stimulus of 0 (no stimulus). These stimulus values were decided based on the experimental setup in the studies we used for validation and calibration (see Section II.2).
II.1.6 Input data
The model does not use input data to represent time-varying processes. However, the model could be extended to allow a time-varying stimulus based on time series input.
II.1.7 Submodels
The model has 2 submodels: (1) the Get_memory submodel which details the agent’s memory process and (2) the Choose_action submodel which describes the mechanics of the agent’s decision-making process.
Get_memory submodel. We assume that the agent’s past actions and stimuli have an exponentially decaying influence on the agent’s memory. If we consider some previous time step $i=t-\Delta t$, the influence $w_{i}$ of the experience at the previous time step $i$ to the memory $m_{t}$ at the current time step $t$ is expressed as:
$w_{i}=\begin{cases}s_{i}e^{-\Delta t/\tau}&\text{if }a_{i}=a_{t}\\ -s_{i}e^{-\Delta t/\tau}&\text{if }a_{i}\neq a_{t}\end{cases}$ (1)
where $a_{t}$ is the agent’s action at the current time step $t$, $a_{i}$ is the agent’s action at the previous time step $i$, $s_{i}$ is the corresponding stimulus that the agent received for action $a_{i}$, $\Delta t$ is the elapsed time between time steps $t$ and $i$, and $\tau$ is the timescale of memory. A higher $\tau$ indicates longer memory, while a smaller $\tau$ means shorter memory. The memory $m_{t}$ of all past actions and stimuli is then computed as:
$m_{t}=\sum_{i=1}^{t-1}{w_{i}}$ (2)
We excluded the current experience at time step $t$ from the memory calculation to enable the current stimulus $s_{t}$ to be weighed differently from the memory $m_{t}$ of past actions and stimuli. Moreover, note that the sign of the influence $w$ in Equation 1 is relative to the current action $a_{t}$. Hence, positive values influence the agent to stay with the current action $a_{t}$ while negative values influence the agent to change the current action $a_{t}$ in the next time step $t+1$.
Choose_action submodel. The agent chooses its next action $a_{t+1}$ at the next time step $t+1$ based on the utility of its current action $a_{t}$ at the current time step $t$. The agent receives a stimulus $s_{t}$ from the environment corresponding to its current action $a_{t}$. The agent also computes a memory value $m_{t}$ for the current action $a_{t}$ using the Get_memory submodel. This memory value indicates the agent’s preference for the current action $a_{t}$ based on its past experiences. The utility $u_{t}$ of the current action $a_{t}$ is then defined as the sum of the environmental stimulus $s_{t}$ and the agent’s internal memory value $m_{t}$:
$u_{t}=s_{t}+m_{t}$ (3)
This utility represents the combination of all factors that the agent considers when making a decision. The utility $u_{t}$ is then passed to a logistic decision function:
$p_{a_{t+1}}=\frac{1}{1+e^{-ku_{t}}}$ (4)
where the steepness $k$ indicates the agent’s certainty on the informational value of any given utility. The logistic decision function returns a probability $p_{a_{t+1}}$ indicating the agent’s likelihood of staying with the current action $a_{t}$ in the next time step $t+1$. Hence, the agent’s next action $a_{t+1}$ will be the same as the current action $a_{t}$ with probability $p_{a_{t+1}}$ or will be the other action with probability $1-p_{a_{t+1}}$.
II.2 Reversal learning experiments
We examined the agent’s behaviour using a simulation of single and serial reversal learning experiments. We followed the experimental setup in Buechel2018 ; Boussard2020 and used their empirical data to calibrate and validate our model. In the single reversal learning experiment on guppies in Buechel2018 , there were two trial periods with 30 trials in the first trial period and 66 trials in the second trial period. Here, we set the second trial period to 30 for uniformity as done similarly in Boussard2020 . We note that we used their empirical data on the large-brained guppies only to make a 1-to-1 comparison with the results they presented in the paper. At every trial, the agent (or the guppy) was given two discs to choose from. During the first trial period, one of the discs had a food reward underneath while the other had none, but during the second trial period, the reward associated with each disc was reversed. We simulated this empirical setup by giving the agent in our model two actions to choose from: A and B. For the first trial period, we set the stimulus $s$ to +1 for action A, signifying the presence of the food reward, and 0 for action B, indicating the absence of the food reward. We then switched the stimulus associated with each action after the reversal point to mimic the reward reversal. The model dynamics of one simulation run of this experiment are illustrated in Figure 3. We also simulated a serial reversal learning experiment which extends the single reversal to a series of reversals as conducted in Boussard2020 . The configurations of the serial reversal learning experiment were similar to the single reversal learning but instead of having only two trial periods separated by one reversal, there were 11 trial periods with ten reversals. These agent-based model parameters are summarized in Table 1. Similar to Buechel2018 ; Boussard2020 , we report the success rate or the fraction of agents that successfully chose the rewarded disc at every trial.
Figure 3: Model dynamics of one simulation run of one agent with a memory decay $\tau=5$ and a logistic function steepness $k=0.5$. (a) The agent selects an action $a_{t}$ (either AA or $B$) at every time step $t$. Highlighted in grey are instances when the agent switches actions. (b) The agent initially receives a stimulus $s_{t}=+1$ for action AA and a stimulus $s_{t}=0$ for action $B$ before the reversal point, but this reward association switches after the reversal. (c) The agent has a memory $m_{t}$ indicating the agent’s preference for each action over time. (d) The agent chooses action $A$ with probability $p_{A}$ based on the stimulus $s_{t}$ and memory $m_{t}$.
II.3 Extended model for serial reversal learning
The result of the serial reversal learning experiments in Boussard2020 revealed a notable characteristic in the agents’ behaviour wherein the agents’ success rate at the end of each trial period decreases over reversals. In our initial simulation runs of the serial reversal learning, we found that our simple two-parameter model was inadequate in capturing this observed phenomenon. Hence, we introduced a third parameter $\theta$ that controls the decay per reversal of the steepness $k$ of the agent’s decision function. Specifically, we decreased the value of $k$ by a factor of $\theta\in[0,1]$ after every reversal period $j$:
$k_{j+1}=k_{j}\times(1-\theta).$ (5)
We can interpret this decay as reducing the impact of information on the agent’s decision by decreasing the steepness of its decision function, which means that higher values of utility $u$ are needed to produce the same probability $p_{a}$.
II.4 Calibration: Approximate Bayesian computation
We used approximate Bayesian computation (ABC) to calibrate our model to empirical data. ABC rejection sampling is a method for estimating the model parameters’ posterior distributions using repeated simulations. Initially, the range of values that each parameter can have is assumed to follow some prior distribution. An ensemble of $N$ models with the parameter values sampled from those priors is then run and the summary statistics of the ensemble are compared with those of the empirical data based on a distance metric. The sampled parameter values are accepted if the distance between the two summary statistics is within the prespecified error threshold $\delta$; otherwise, they are rejected. New parameter values are sampled from the prior distributions and the process repeats until $l$ samples are accepted. These accepted samples characterise the posterior distributions of the model parameters which can be utilised to simulate the observed behaviours from the empirical data. Here, we assumed a uniform prior distribution for each parameter, used the success rate over trials as the summary statistics, and chose the mean squared error as the distance metric.
      Symbol Description Value
(Single reversal)
Value
(Serial reversal)
     Agent-Based Model
      $n$ Number of trial periods 2 11
      $r$ Number of trials per trial period 30 30
      $\tau$ Memory decay See Figure 4b See Figure 5b
      $k$ Steepness of the logistic decision function See Figure 4b See Figure 5b
      $\theta$ Decay factor applied to $k$ per trial period 0 See Figure 5d
     Approximate Bayesian Computation
      NN Ensemble count 100 100
      $\delta$ Error threshold 0.01 0.01
      $l$ Sample count 1,000 1,000
      $U(\tau_{min},\tau_{max})$ Prior distribution of $\tau$ $U(0,50)$ $U(0,50)$
      $U(k_{min},k_{max})$ Prior distribution of $k$ $U(0,1)$ $U(0,1)$
      $U(\theta_{min},\theta_{max})$ Prior distribution of $\theta$ $U(0,0)$ $U(0,1)$
Table 1: Parameters of the agent-based model and the approximate Bayesian computation used for the single and serial reversal learning simulations.
For the serial reversal learning simulations, the calibration was performed in two stages. The first stage of the calibration process is identical to the one we used for the single reversal learning simulations where the posterior distributions of the memory decay $\tau$ and the steepness $k$ are calculated for the first two trial periods (separated by a single reversal). In the second stage, we computed the posterior distribution of $\theta$ that captures the behaviour across multiple reversals, with the $\tau$ and $k$ parameter values fixed at the maximum a posteriori (MAP) parameter estimate obtained from the first stage. We note that this two-stage calibration procedure is not guaranteed to converge to a maximum likelihood posterior over the joint distribution of $k$, $\tau$, and $\theta$, but nevertheless performs well enough to allow adequate reproduction of the observations after calibration. The parameter ranges we used in our ABC methods are listed in Table 1.
III Results
Here, we present the posterior distributions of the model parameters obtained from the calibration process using ABC and compare the simulated behaviours given those parameters with real behaviours observed from empirical reversal learning experiments.
III.1 Single Reversal Learning
The simulations recovered the observed trajectory of the success rate given parameter values drawn from the posterior distributions of the memory decay $\tau$ and steepness $k$ obtained from the ABC as shown in Figure 4. The success rate shows the average success across 100 agents with the same $\tau=9.7$ and $k=0.26$ parameters. At the start of the first trial period, the success rate is only about 50% because the agents are still unaware of the task at hand. The agents eventually learn which option has the reward after about 20 trials wherein the success rate reaches 80%, which is the criterion set in Buechel2018 to indicate that an agent learned to discriminate the rewarded and non-rewarded options. When the reward is reversed after the first trial period, the success rate drops from about 90% to 10%. This drop in performance indicates the effect of memory on the agents’ decision-making process. Despite the change in the stimulus from the environment, the agents’ memory still favours the previously rewarded option. The agents eventually learn the new reward association over time such that the success rate steadily increases back to 80%.
Figure 4: Model calibration to single reversal learning experiments. The trajectory in (a) shows the success rate over an ensemble of 100 agents in a single reversal learning experiment in which the reward reversal occurs after 30 trials. The black trace is the experimental data from Buechel2018 , the grey traces are the simulation results using all accepted ABC samples, and the orange trace is the simulation result using the MAP parameter estimate. The ABC posterior is shown in (b) with the MAP parameter estimate indicated by the orange point.
III.2 Serial Reversal Learning
As shown in Figure 5a, the first calibration stage of the serial reversal learning simulations also recovered the observed success rate in the first of ten reversals from the experimental data from Boussard2020 . The agents’ success rate goes beyond 80% after several trials, exhibits a huge drop right after the reversal point, and then rises to about 90% eventually. With $\theta=0.12$ obtained from the posterior distribution of $\theta$ after the second calibration stage, the simulated data captures the decline in success rate in the succeeding reversals as depicted in Figure 5c. The success rate after the first reversal goes back up to about 90% but only reaches about 70% after the tenth reversal.
Figure 5: Two-stage model calibration to serial reversal learning experiments. The trajectories in (a) and (c) show the success rate over an ensemble of 100 agents in the first reversal after the first calibration stage and in all ten reversals of a serial reversal learning experiment after the second calibration stage, respectively. The black traces are the experimental data from Boussard2020 , the grey traces are the simulation results using all accepted ABC samples, and the orange traces are the simulation results using the MAP parameter estimate. The ABC posterior is shown in (b) for the first calibration stage and (d) for the second calibration stage with the MAP parameter estimate indicated by the orange point.
IV Discussion
Representation of individual behaviour change in computational models of real-world systems is essential to understanding and predicting phenomena emerging from collective behaviour. However, many behavioural change models rely on arbitrary assumptions unsupported by quantitative empirical data which may limit the validity of simulation results. Here, we demonstrated a simple model of behaviour change that can simulate actual behaviours observed in single reversal learning experiments. We further showed that the model can be easily extended to simulate more complex behaviours in serial reversal learning experiments with longer time scales.
We developed an agent-based model wherein an individual’s behaviour is characterised by three parameters: (1) the memory decay $\tau$ that indicates the extent to which an individual uses previous experiences to make decisions, (2) the steepness $k$ of the logistic decision function that determines an individual’s certainty on the value of given information, and (3) the factor $\theta$ that specifies the decay of an agents certainty about information value over multiple reward reversals. We calibrated these model parameters using ABC given the empirical data from Buechel2018 and Boussard2020 . As shown in Figures 4 and 5, the simulated behaviours using those identified parameters are well-matched to the ensemble statistics observed in the empirical experiments.
These results indicate that our proposed behavioural change model satisfies mechanistic adequacy, tangibility, and tractability. First, with only two parameters, our model was able to simulate the following qualitative patterns observable in reversal learning data: (1) the capacity to learn associations between decisions and rewards, (2) stochasticity in the decision process, and (3) transient persistence of learned behaviours after the original stimulus is reversed. Further, by adding one additional parameter, our model reproduced the long-term behaviour observed in experiments with multiple reversals. In all cases, calibration to observed behaviour located compact regions in the parameter spaces of our models, indicating that the parameters can be efficiently identified from empirical data. Our work thus demonstrates the feasibility of modelling and simulating real-world behaviours using behavioural data.
Individual behaviours in reversal learning experiments demonstrate a balance between behavioural flexibility (adaptation in response to changing stimuli) and persistence (the tendency for learned behaviour to persist even when rewards are inconsistent). Such a balance is essential for individuals to survive in a dynamically changing world, so it is not surprising that the observed behaviour in guppies in the single reversal learning experiment examined here is consistent with observations of bees Dyer2014 , birds Rayburn-reeves2011 , rodents Le2023 , monkeys Hassett2017 , and humans Rayburn-reeves2011 ; Xia2021 . The decline in success rate observed in the serial reversal experiments we studied suggests a complex mechanism of information processing in which the perceived value of information appears to decay after many fluctuations in the reward signal. We speculate this may be analogous to behavioural phenomena in humans such as declining adherence to protective behaviours during the latter stages of the COVID-19 pandemic Petherick2021 . Designing computational models to have the mechanistic adequacy to simulate such general and specific behavioural change processes could, for example, help better understand animal foraging behaviour Dyer2014 or enable the design of more effective policy interventions during infectious disease outbreaks Petherick2021 . However, to be useful, such models need to have tractable computational complexity and cost, while also maintaining tangible sets of observable inputs and outputs.
In order to calibrate and validate the model to reproduce such real-world behaviours, there is a need for suitable behavioural data. However, because those data are difficult to collect, many existing models rely on assumptions without any empirical support. Moreover, some models that incorporate data rely only on behavioural surveys Durham2012 ; Fierro2013 ; Zhong2013 ; Bayham2015 ; Karimi2015 , which are difficult to scale to larger populations and often characterise individuals’ behavioural intentions that are not necessarily consistent with their actual behaviours Verelst2016 ; Weston2020 . Behavioural data capturing individuals’ actual actions provides a more accurate representation of real-world behaviours. For example, mobility data such as GPS trajectories from mobile phones can be readily collected under different scenarios and conditions Badr2020 ; Xiong2020 ; Ilin2021 . Other indicators of behaviour such as television viewing duration Springborn2015 or home internet usage Zachreson2021 can help quantify ensembles of discrete individual choices. Developing models of individual decision making that can be calibrated to such ensemble data is important for ensuring that simulated behaviours are tangible and valid in the context of the collective behaviour of interest such as the endogenous response to infectious disease outbreaks.
Beyond reproducing real-world behaviours, ensuring that the model parameters that generated such behaviours are identifiable from empirical data is also crucial when designing behavioural change models. Because models are often used to inform decision making, it is desirable to have full and explicit information on what drives a model’s dynamics to improve its interpretability and reliability. By using ABC to calibrate the model, we can characterise the posterior distributions of the model parameters and demonstrate the uncertainty in the model’s dynamics and predictions. Quantifying this model uncertainty could make the analysis and utilisation of models more robust and effective because it provides information on how to interpret their results and on which instances they should be relied on Edeling2021 .
Previous reviews of behavioural change models have also highlighted that many models are not informed by behavioural change theories which provide foundations for understanding, predicting, and influencing individual behaviours. Despite such theories being used in designing and implementing nationwide public health policies and disaster preparedness and response strategies Michie2011 ; Angus2013 ; West2020 , they are rarely incorporated into models that are also developed to support decision making in those scenarios Weston2020 . We designed behaviour in our agent-based model based broadly on the COM-B model for behaviour change which states that capability (C), opportunity (O), and motivation (M) are the key factors that influence behaviour (B) Michie2011 . In our model, the agent’s decision function and internal memory could represent its capability and motivation to behave in a certain way, while the external stimulus could symbolise the opportunity provided by the environment to make the behaviour possible. Our proposed model is a simple computational representation consistent with the COM-B model which can be extended to account for the many other components in each of those factors.
Our study has several limitations. The first is that we only considered a binary set of actions, which many decision-making processes can be reduced to. However, the model can be extended to incorporate more than two actions using other decision functions like a softmax function Sutton2018 . The second limitation is that we used an aggregate behavioural measure to calibrate the model parameters. While those parameters may not be characteristic of each individual’s behaviour, our approach could be extended to represent ensembles of agents with values of individual-level model parameters drawn from population-level hyperdistributions. In practical terms, the approach presented here can be viewed as a technique for building individual-level models that are consistent with ensemble observations. This approach is consistent with typical observables of real-world systems, which are often quantified in aggregate for subpopulations such as communities or demographic strata Springborn2015 ; Badr2020 ; Xiong2020 ; Ilin2021 .
V Conclusion
In summary, we developed a simple yet extensible agent-based model of behaviour change that is mechanistically adequate to simulate reversal learning. We calibrated the parameters of our model from empirical behavioural data using approximate Bayesian computation and validated that the model reproduces behaviours observed in single and serial reversal learning experiments given those parameters. This work is a step towards developing computational models of behaviour change whose parameters are fully identifiable from actual behavioural data and are grounded in psychological theories. Such models are essential for understanding and predicting how behavioural changes affect emergent real-world phenomena with better accuracy, reliability, and practicality.